Consider the reaction.

PCl5(g)−⇀PCl3(g)+Cl2(g)Kc=0.0420

The concentrations of the products at equilibrium are [PCl3]=0.170 M and [Cl2]=0.190 M . What is the concentration of the reactant, PCl5 , at equilibrium?
1
Expert's answer
2020-03-30T10:48:38-0400

Equilibrium concentration of PCl3= 0.170 M

Equilibrium concentration of Cl2= 0.190 M

The equilibrium constant for the reaction Kc= 0.0420

The equation relating the equilibrium constant to the equilibrium concentrations of the reactants and products is:


"Kc=\frac{PCl3\times Cl2}{PCl5}"


"0.0420=\frac{0.170\times0.190}{PCl5}"


"PCl5=\frac{0.0420}{0.0323}"


PCl5=1.3000